Ncert Solutions Maths class 12th
Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th
- 2.5k
Questions
- 0
Discussions
- 0
Active Users
- 65
Followers
Have a question about Ncert Solutions Maths class 12th? Ask our experts
Keep it short & simple. Type complete word. Avoid abusive language.
If the foot of the perpendicular drawn from the point on a line passing through is , then is equal to
If the foot of the perpendicular drawn from the point on a line passing through is , then is equal to
(5/3, 7/3, 17/3)
AD ⋅ PD = 0
((5/3-α)i?+(7/3-7)j?+(17/3-1)k?) . (2/3i?+7/3j?+8/3k?) = 0
(5/3-α)(2/3) + (-14/3)(7/3) + (14/3)(8/3) = 0
A(α, 7, 1) D(7/3, 7/3, 12/3)
⇒ 3α = 12
α = 4
If the system of linear equations,
<!--[endif]--><!--[endif]-->
<!--[endif]--><!--[endif]-->
<!--[endif]--><!--[endif]-->
has more than two solutions, then is equal to
If the system of linear equations,
<!--[endif]--><!--[endif]-->
<!--[endif]--><!--[endif]-->
<!--[endif]--><!--[endif]-->
has more than two solutions, then is equal to
Δ = |1 1 1| = 0
|1 2 3|
|3 2 λ|
⇒ 1(2λ - 6) - 1(λ - 9) + 1(-4) = 0
⇒ λ = 1
Δx = |6 1 1|
|10 2 3| = 0
|μ 2 λ|
⇒ 2λ + μ = 16
⇒ μ = 14
μ - λ² = 14 - 1 = 13
Let the vectors a,b,c be such that |a|=2, |b|=4 and |c|=4. If the projection of b on a is equal to the projection of c on a and b is perpendicular to c, then the value of |a+b-c| is
Let the vectors a,b,c be such that |a|=2, |b|=4 and |c|=4. If the projection of b on a is equal to the projection of c on a and b is perpendicular to c, then the value of |a+b-c| is
6.00
b·a = c·a
|a+b-c|² = |a|²+|b|²+|c|²+2(a·b - b·c - a·c)
= 4+16+16+2(a·b - 0 - a·b) = 36
⇒ |a+b-c| = 6
In a workshop, there are five machines and the probability of any one of them to be out of service on a day is <!-- [if gte msEquation 12]> 1 4
<!--[endif]--><!--[endif]-->. If the probability that at most two machines will be out of service on the same day is , then is equal to:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
In a workshop, there are five machines and the probability of any one of them to be out of service on a day is <!-- [if gte msEquation 12]>
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
Probability (at most two machines will be out of service) = (3/4)³ . k
⇒ ?C?(1/4)?(3/4)? + ?C?(1/4)¹(3/4)? + ?C?(1/4)²(3/4)³ = (3/4)³ . k
⇒ 17/8 (3/4)³ = (3/4)³ . k
⇒ k = 17/8
Let and be two independent events such that and . Then, which of the following is TRUE?
Let and be two independent events such that and . Then, which of the following is TRUE?
A and B are independent events so P(A/B) = 1/3
If for some α ∈ R, the lines L₁: (x+1)/2 = (y-2)/(-1) = (z-1)/1 and L₂: (x+2)/α = (y+1)/(5-α) = (z+1)/1 are coplanar, then the line L₂ passes through the point:
If for some α ∈ R, the lines L₁: (x+1)/2 = (y-2)/(-1) = (z-1)/1 and L₂: (x+2)/α = (y+1)/(5-α) = (z+1)/1 are coplanar, then the line L₂ passes through the point:
Line are coplanar
so | [α, 5-α, 1], [2, -1, 1], | = 0
−5α + (α – 5)3 + 7 = 0
-2α = 8 ⇒ α = −4
⇒ L? : (x+2)/-4 = (y+1)/9 = (z+1)/1
Now by cross checking option (A) is correct.
w = I + j - 3k
p = 3i - j + k
q = -3i + 2j + 4k
p × q = |i, j, k; 3, -1, 1; -3, 2, 4| = -6i - 15j + 3k
S.D. = |AD. (p × q)|/|p × q| = | (36+225+9)|/√ (36+225+9) = 3√30
Let be a function of satisfying where is a constant and . Then at , is equal to:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
Let be a function of satisfying where is a constant and . Then at , is equal to:
<!-- [if !supportLineBreakNewLine]-->
<!--[endif]-->
Differentiating
y.
Put and
Let a curve y = f(x) pass through the point (2, (logₑ2)²) and have slope 2y/(xlogₑx) for all positive real value of x. Then the value of f(e) is equal to……….
Let a curve y = f(x) pass through the point (2, (logₑ2)²) and have slope 2y/(xlogₑx) for all positive real value of x. Then the value of f(e) is equal to……….
dy/dx = 2y/ (xlnx).
dy/y = 2dx/ (xlnx).
ln|y| = 2ln|lnx| + C.
ln|y| = ln (lnx)²) + C.
y = A (lnx)².
(ln2)² = A (ln2)². ⇒ A=1.
y = f (x) = (lnx)².
f (e) = (lne)² = 1² = 1.
A fair coin is tossed n-times such that the probability of getting at least one head is at least 0.9. Then the minimum value of n is………….
A fair coin is tossed n-times such that the probability of getting at least one head is at least 0.9. Then the minimum value of n is………….
P (at least one head) = 1 - P (no heads) = 1 - (1/2)? ≥ 0.9.
0.1 ≥ (1/2)?
10 ≤ 2?
n=3, 2³=8. n=4, 2? =16.
Minimum value of n is 4.